# How much can we forget about Data Contamination?

### v1

@@ Abstract
The leakage of benchmark data into the training data has emerged as a significant challenge for evaluating the capabilities of large language models (LLMs). In this work, we use experimental evidence and theoretical estimates to challenge the common assumption that small-scale contamination renders benchmark evaluations invalid. First, we experimentally quantify the magnitude of benchmark overfitting based on scaling along three dimensions: The number of model parameters (up to 1.6B), the number of times an example is seen (up to 144), and the number of training tokens (up to 40B). We find that if model and data follow the Chinchilla scaling laws, minor contamination indeed leads to overfitting. At the same time, even 144 times of contamination can be forgotten if the training data is scaled beyond five times Chinchilla, a regime characteristic of many modern LLMs. We then derive a simple theory of example forgetting via cumulative weight decay. It allows us to bound the number of gradient steps required to forget past data for any training run where we know the hyperparameters of AdamW. This indicates that many LLMs, including Llama 3, have forgotten the data seen at the beginning of training. Experimentally, we demonstrate that forgetting occurs faster than what is predicted by our bounds. Taken together, our results suggest that moderate amounts of contamination can be forgotten at the end of realistically scaled training runs.
@@ 1 Introduction
Our investigation reveals that the natural forgetting dynamics of gradient descent [66, 39] is the reason why increasing the number of tokens alleviates the impact of contamination. Concretely, we show that training on five times Chinchilla [35] of clean data can cause a model to forget even 144 times repeated training examples (Section 4.2). Forgetting the bulk of the impact of a training example can occur rapidly, a point that we demonstrate using OLMo-1B [30] (Section 4.3). What is the reason for this rapid forgetting? We show that exposure to novel data is important. Interestingly, models tend to exhibit the strongest overfitting on examples seen repeatedly throughout training, even compared to those seen during the end (Section 4.2).
Because running pre-training experiments is expensive, we also ask to what degree forgetting can be explained by the training dynamics of gradient descent. We show that the weight decay parameter and learning rate schedule of the AdamW optimizer [49] play a key part in forgetting past training examples (Section 5). Concretely, we derive a simple theory of forgetting via cumulative weight decay (Section 5.1) and show that it provides an upper bound on empirical forgetting, which usually occurs faster (Section 5.2). The key point is that this approach allows us to gauge the degree of forgetting present in any training run for which the optimization hyperparameters are known. It even allows us to approximate how the final model parameters balance the gradient updates from different stages of training. Overall, our analysis indicates that many LLMs, including OLMo-7B and Llama 3 405B [22], have forgotten the data seen at the beginning of their training run.
@@ 2 Related Work
Forgetting. In machine learning, the term forgetting is frequently associated with “catastrophic” forgetting, where learning new tasks hurt the performance at previously solved tasks [48]. In the context of LLMs, catastrophic forgetting can occur during fine-tuning [51] or continual learning [36]. In contrast, this paper studies forgetting as a potential “natural” phenomenon of learning [67]. Tirumala et al. [66] study forgetting in language modeling and find, similar to Toneva et al. [67], that forgetting can be exponentially slow. In contrast, Jagielski et al. [39] find that models empirically do forget examples over time. In concurrent work, Pagliardini et al. [58] propose to add a second momentum term to the AdamW optimizer, and show that this slows down the forgetting of past gradients.
Data Attribution. Data attribution methods [42, 38, 59] aim to identify data points responsible for specific model behaviors. We ask how much a model’s benchmark performance is influenced by seeing the example during training, which broadly falls within this field [31, 16]. Importantly, we directly measure the influence of contaminated examples through retraining, avoiding the approximation errors that can occur when using data attribution methods [42, 26] for large-scale models [5, 4]
@@ 3 Background and Methods
Models and Training Data. We train language models of up to 1.6B parameters using the architecture and hyperparameters from the GPT-3 paper [10]. For this, we adopt the llm.c codebase. The training data is the 100BT split of the FineWeb-Edu dataset [50]. We also train OLMo-1B [30] using the corresponding code and data [65].
@@ 3.1 We consider the regime of n-times Chinchilla
3.2 We evaluate on a mix of seven different benchmarks
3.3 We filter near-duplicate Benchmark Questions
Our method requires that there are no side effects from contaminating the training data with one question on the evaluation of another question. However, upon closer inspection, it turns out that all the commonly used benchmarks from the literature contain questions that are either near-duplicates or where the context of one question contains the answer to another question (for example, because the same text document was used to create multiple questions). This is illustrated in Figure 1, which depicts two near-duplicate questions on HellaSwag and questions from ARC-Easy and MMLU that are cross-benchmark duplicates. To address this problem, we perform extensive filtering, removing duplicate questions where the length-normalized Levenshtein distance falls below a certain threshold [43, 56]. This is documented in Supplement A.3, where we also describe an experiment to verify that our method is not invalidated by near-duplicate questions.
4 Experimental Results
We now present our main experimental results. We being in Section 4.1 by discussing the scaling in model parameters, training tokens, and repetitions in the training data. The following Section 4.2 discusses various experiments on forgetting. The first two sections rely on training small GPT-3 models. Section 4.3 complements this with an analysis of OLMo-1B [30].
4.1 Contamination scales with Model, Data, and Repetitions
@@
Table 1: Accuracy of the Chinchilla models.
@@ 4.2 Contamination can be completely Forgotten
Figure 3: The natural forgetting dynamic of neural network training. (a) The development of the cross-entropy loss difference between contaminated and holdout benchmark questions over the course of training. Contamination occurs between the first and second Chinchilla (1 and 2 on the x-axis). (b) Accuracy gaps after training for 3 Chinchilla. (c) Accuracy gaps after training for 7 Chinchilla. (d) Same as (a). (e)+(f) The accuracy gap depends on the average position of an example in the training data. Mean and bootstrapped 90% confidence intervals.
In the previous Section 4.1, we saw that the accuracy gap due to contamination decreases in the number of tokens up to the point where even 12 repetitions of a benchmark question in the training data can become insignificant. In this Section, we identify the natural forgetting dynamic of neural network training as the reason for this effect. We discuss how quickly forgetting occurs, whether examples are completely forgotten, and what kind of repetition makes a model remember.
4.3 Contamination and rapid forgetting in OLMo-1B
@@
Figure 4: Contamination and forgetting in OLMo-1B. We contaminate the OLMo-1B checkpoint at gradient step 369,000 four times with different benchmarks. This causes an average accuracy increase of 15 percentage points. We then continue pre-training for 0.5% of the remaining training time, leading to a reduction of 84% of the accuracy increase due to contamination. In this Figure, different colors simply correspond to different benchmarks, and the grey line depicts the clean accuracy without contamination. Mean and bootstrapped 90% confidence intervals.
@@
In the previous sections, we trained small GPT-3 models from scratch. In this Section, we complement this analysis by pre-training from an intermediate OLMo-1B checkpoint [30]. Similar to the analysis in Section 4.2, we insert the benchmark data at a specific point into the training data and then measure the subsequent forgetting. Unlike in the previous Section, we now insert the entire benchmark data – we already have a “clean” baseline from the original OLMo-1B training run. We insert each benchmark question four times and contaminate with four different benchmarks: HellaSwag [72], WinoGrande 62, ARC-Easy [19], and PiQA [7].
@@
Figure 4 depicts the result of the experiment. The effect of contamination is visible from the five leftmost points of every plot. The leftmost point corresponds to the uncontaminated model, and the next four points each depict the effect of one time contamination. Again, we see that the immediate effect of contamination is significant, leading to an average accuracy increase of 15 percentage points across the different benchmarks. At the same time, we also see that the effect of contamination decays considerably as we continue training. To contextualize this result, note that Figure 4 depicts less than 2000 gradient steps. The pre-training stage of OLMo-1B model consists of 739,328 gradient steps. This means that Figure 4 depicts less than 1% of the total forgetting until pre-training is done.
5 What is the role of weight decay in forgetting?
@@
Figure 5: Theoretical estimates of forgetting and approximate model weight composition at the end of training. Top Row: The cumulative weight decay $w_{t_{1}}^{t_{2}}$ as defined in equation (3) for different training runs. The figures depict the decay of the gradient updates for every decile of the training run, indicated in different colors. Bottom Row: The approximate composition of the final model weights in terms of the gradient updates from different deciles of the training run. The deciles are indicated in colors depicted in the legend below the plot.
In the previous Section 4, we have seen that forgetting is an important empirical property of LLM training. In this Section, we show that the weight decay parameter and learning rate schedule of the AdamW optimizer play a key part in forgetting past training examples. This offers a novel perspective on the interplay between these two parameters, usually seen in terms of generalization and training stability [69, 74, 44, 3].
5.1 Weight Decay as a Mechanism for Forgetting Training Examples
Consider the parameter update of AdamW at gradient step $t\geq 1$. It consists of two decoupled updates [60, 61]: A weight decay update given by
$\hat{\theta}_{t}=\theta_{t-1}-\gamma\lambda_{t}\theta_{t-1},$ (1)
and a gradient update given by
$\theta_{t}=\hat{\theta}_{t}-\lambda_{i}\,\hat{m_{t}}/(\sqrt{\hat{v_{t}}+\epsilon}).$ (2)
Here, $\theta_{t}$ are the model parameters, $\lambda_{t}$ is the learning rate, $\gamma$ is the weight decay parameter, and $\hat{m_{t}}$ and $\hat{v_{t}}$ are first- and second-order moment estimates of the gradient. Denoting the model weights at initialization by $\theta_{0}$, and the adaptive gradient by $\hat{g}_{t}=\hat{m_{t}}/(\sqrt{\hat{v_{t}}+\epsilon})$, we can iterate (1) and (2) to obtain
$\theta_{T}=w_{0}^{T}\,\theta_{0}+\sum_{t=1}^{T}w_{t}^{T}\,\lambda_{t}\,\hat{g}_{t}\quad\quad\text{where}\quad w_{t_{1}}^{t_{2}}=\prod_{i=t_{1}}^{t_{2}}(1-\lambda_{i}\gamma).$ (3)
Here, the weights $w_{t_{1}}^{t_{2}}$ account for the cumulative weight decay between gradient step $t_{1}$ and $t_{2}$. Intuitively, the model weights after $t$ gradient steps are a weighted average of the initial model weights and all the adaptive gradient updates up to time step $t$. This is not specific to the AdamW optimizer and applies to every optimizer with weight decay. Analyzing equation (3) reveals a critical factor influencing forgetting: The exponential decay of the $w^{t_{2}}_{t_{1}}$ with respect to increasing gap $t_{2}-t_{1}$ in the optimization steps. In other words, the longer an update occurs in the past (i.e., the larger $t_{2}-t_{1}$), the more the contribution of the update $\hat{g}_{t_{1}}$ is being scaled down due to the exponential decay of weights $w^{t_{2}}_{t_{1}}$. We can describe the evolution of these weights as the function of the time $T=t_{2}-t_{1}$.
Proposition 1.
(The Decay of Past Gradients) The number of optimization steps $T=t_{2}-t_{1}$ that are required to make the contribution of a model update at time $t_{1}$ small, that is $w^{t_{2}}_{t_{1}}\leq\epsilon$ for some small ϵ+italic-$\epsilon\in\mathbb{R}^{+}$, scales as $T\gtrsim\frac{log(1/\epsilon)}{\gamma\lambda_{\text{avg}}}$, where $\lambda_{\text{avg}}=\frac{1}{T}\sum_{t={t_{1}}}^{t_{2}}\lambda_{t}$ is the average learning rate of the optimizer between $t_{1}$ and $t_{2}$.
We present a proof in the Supplement A.5. If $w_{t_{2}}^{t_{1}}\leq\epsilon$ for some sufficiently small $\epsilon$, the term $w_{t_{2}}^{t_{1}}\lambda_{t_{1}}\hat{g}_{t_{1}}$ vanishes from the sum in (3). Intuitively, this is the same as saying that the gradient update at time step $t_{1}$ has been forgotten.

### v3

@@ Abstract
The leakage of benchmark data into the training data has emerged as a significant challenge for evaluating the capabilities of large language models (LLMs). In this work, we challenge the common assumption that small-scale contamination renders benchmark evaluations invalid. First, we experimentally quantify the magnitude of benchmark overfitting based on scaling along three dimensions: The number of model parameters (up to 1.6B), the number of times an example is seen (up to 144), and the number of training tokens (up to 40B). If model and data follow the Chinchilla scaling laws, minor contamination indeed leads to overfitting. At the same time, even 144 times of contamination can be forgotten if the training data is scaled beyond five times Chinchilla, a regime characteristic of many modern LLMs. Continual pre-training of OLMo-7B corroborates these results. Next, we study the impact of the weight decay parameter on example forgetting, showing that empirical forgetting occurs faster than the cumulative weight decay. This allows us to gauge the degree of forgetting in large-scale training runs, indicating that many LLMs, including Lllama 3 405B, have forgotten the data seen at the beginning of training.
@@ 1 Introduction
Our investigation reveals that the forgetting dynamics of gradient descent [66, 39] is the reason why increasing the number of tokens alleviates the impact of contamination. Concretely, we show that training on five times Chinchilla [35] of clean data can cause a model to forget even 144 times repeated training examples (Section 4.2). Forgetting also occurs for larger models, a point that we demonstrate using OLMo-7B [30] (Section 4.3). What is the reason for the forgetting? We show that exposure to novel data is important. Interestingly, models tend to exhibit the strongest overfitting on examples seen repeatedly throughout training, even compared to those seen during the end (Section 4.2).
Because running pre-training experiments is expensive, we also ask to what degree forgetting can be explained by the training dynamics of gradient descent. We find that the weight decay parameter and learning rate schedule of the AdamW optimizer [49] play a key part in forgetting past training examples (Section 5). Concretely, we demonstrate that the cumulative weight decay (Section 5.1) bounds empirical forgetting (Section 5.2). This approach allows us to gauge the degree of forgetting in large-scale training runs by analyzing the optimization hyperparameters (Section 5.3). It also allows us to approximate how the final model parameters balance the gradient updates from different stages of training. Our analysis indicates that many LLMs, including OLMo-7B and Llama 3 405B [22], have forgotten the data seen at the beginning of training.
@@ 2 Related Work
Forgetting. In machine learning, the term forgetting is frequently associated with “catastrophic” forgetting, where learning new tasks hurt the performance at previously solved tasks [48]. In the context of LLMs, catastrophic forgetting can occur during fine-tuning [51] or continual learning [36]. In contrast, this paper studies forgetting as a potential “natural” phenomenon of learning [67]. Tirumala et al. [66] study forgetting in language modeling and find, similar to Toneva et al. [67], that forgetting can be exponentially slow. In contrast, Jagielski et al. [39] find that models empirically do forget examples over time. Pagliardini et al. [58] propose to add a second momentum term to the AdamW optimizer, and show that this slows down the forgetting of past gradients.
Data Attribution. Data attribution methods [42, 38, 59] aim to identify data points responsible for specific model behaviors. We ask how much a model’s benchmark performance is influenced by seeing the example during training, which broadly falls within this field [31, 16]. Importantly, we directly measure the influence of contaminated examples through explicit retraining, avoiding the approximation errors that can occur when using data attribution methods [42, 26] for large-scale models [5, 4]
@@ 3 Background and Methods
Models and Training Data. We train language models of up to 1.6B parameters using the architecture and hyperparameters from the GPT-3 paper [10]. For this, we adopt the llm.c codebase. The training data is the 100BT split of the FineWeb-Edu dataset [50]. We also train OLMo-7B [30] using the corresponding code and data [65].
@@ 3.1 We Consider the Regime of n-times Chinchilla
3.2 We Evaluate on a Mix of Different Benchmarks
3.3 We Filter Near-Duplicate Benchmark Questions
Our method requires no side effects from contaminating the training data with one question on the evaluation of another question. However, upon closer inspection, it turns out that all the commonly used benchmarks from the literature contain questions that are either near-duplicates or where the context of one question contains the answer to another question (for example, because the same text document was used to create multiple questions). To address this problem, we perform extensive filtering, removing duplicate questions where the length-normalized Levenshtein distance falls below a certain threshold [43, 56]. The issue of duplicate benchmark questions is detailed in Supplement A.3, where we also describe an experiment to verify that near-duplicate questions do not invalidate our method.
4 Experimental Results
We now present our main experimental results. We being in Section 4.1 by discussing the scaling in model parameters, training tokens, and repetitions in the training data. The following Section 4.2 discusses various experiments on forgetting. The first two sections rely on training small GPT-3 models. Section 4.3 complements this with an analysis of OLMo-7B [30].
4.1 Contamination Scales with Model, Data, and Repetitions
@@
Table 1: Accuracies of the Chinchilla-optimal models. The table depicts the absolute accuracies of the Chinchilla-optimal models for holdout and contaminated benchmark questions.
@@ 4.2 Contamination Can be Completely Forgotten
Figure 2: The forgetting dynamic of neural network training. (a) The development of the cross-entropy loss difference between contaminated and holdout benchmark questions over the course of training. Contamination occurs between the first and second Chinchilla (1 and 2 on the x-axis). (b) Accuracy gaps after training for 3 Chinchilla. (c) Accuracy gaps after training for 7 Chinchilla. (d) Same as (a). (e)+(f) The accuracy gap depends on the average position of an example in the training data. Mean and bootstrapped 90% confidence intervals.
In the previous Section 4.1, we saw that the accuracy gap due to contamination decreases in the number of tokens up to the point where even 12 repetitions of a benchmark question in the training data can become insignificant. In this Section, we identify the forgetting dynamic of neural network training as the reason for this effect. We discuss how quickly forgetting occurs, whether examples are completely forgotten, and what kind of repetition makes a model remember.
@@
Figure 3: Contamination and forgetting in OLMo-7B. We contaminate the OLMo-7B checkpoint at gradient step 300,000 four times with different benchmarks. This causes an average accuracy increase of 17 percentage points. We then continue pre-training for 13% of the remaining training time, leading to a reduction of 87% of the accuracy increase due to contamination. In this figure, different colours correspond to different benchmarks, and the grey line depicts the clean accuracy without contamination. Mean and bootstrapped 90% confidence intervals.
4.3 Contamination and Forgetting in OLMo-7B
@@
In the previous sections, we trained small GPT-3 models from scratch. In this Section, we complement this analysis by continual pre-training from an intermediate OLMo-7B checkpoint [30]. Similar to the analysis in Section 4.2, we insert the benchmark data at a specific point into the training data and then measure the subsequent forgetting. Unlike in the previous Section, we now insert the entire benchmark data – we already have a “clean” baseline from the original OLMo-7B training run. We insert each benchmark question four times and contaminate with four different benchmarks: HellaSwag [72], WinoGrande 62, ARC-Easy [19], and PiQA [7].
@@
Figure 3 depicts the result of the experiment. The leftmost point in every plot corresponds to the uncontaminated model, and the following four points depict the effect of contamination. The plots then depict how the effect of contamination is forgotten as we continue training. Similar to the results in the previous sections, the immediate effect of contamination is significant, leading to an average accuracy increase of 17 percentage points across the different benchmarks. At the same time, the effect of contamination decays considerably as we continue training. After continual pre-training for 13% of the remaining training time, the evaluation on WinoGrande and ARC-Easy is no longer significantly different from the uncontaminated evaluation. On HellaSwag and PiQA, the effect of contamination has decayed to approximately 2 percentage points of overfitting.
Supplement Figure 7 depicts the same experiment with OMLo-1B. The behaviour of the 1B parameter model is qualitatively similar to the 7B parameter model seen in this section. However, forgetting for the 1B parameter model occurs faster, corroborating the results from Section 4.1.
5 Weight Decay and Forgetting
(a) Weight Decay 50
(b) Weight Decay 5
(c) Weight Decay 1
(d) Weight Decay 0.1
Figure 4: Empirical forgetting occurs faster than the cumulative weight decay. We continue training the contaminated 124M parameter model from Section 4.2. The figure depicts empirical forgetting (dashed line) and the cumulative weight decay (solid line) for four different choices of the weight decay parameter. Forgetting depends strongly on the weight decay parameter, as is evident from the very different x-axis scales, ranging from 120 gradient steps in (a) to 62500 gradient steps in (d).
In the previous Section 4, we have seen that forgetting is an important empirical property of LLM training. Which factors contribute to the phenomenon? In this section, we show that the weight decay parameter and learning rate schedule of the AdamW optimizer play a part in forgetting past training examples. This offers a novel perspective on the weight decay parameter, usually seen in terms of generalization and training stability [69, 74, 44, 3].
5.1 The Cumulative Weight Decay
Consider the parameter update of AdamW at gradient step $t\geq 1$. It consists of two decoupled updates: A weight decay update given by
$\hat{\theta}_{t}=\theta_{t-1}-\gamma_{t}\lambda\theta_{t-1},$ (1)
and a gradient update given by
$\theta_{t}=\hat{\theta}_{t}-\gamma_{t}\,\hat{m}_{t}/(\sqrt{\hat{v}_{t}}+\epsilon).$ (2)
Here, $\theta_{t}$ are the model parameters, $\gamma_{t}$ is the learning rate, $\lambda$ is the weight decay parameter, and $\hat{m_{t}}$ and $\hat{v_{t}}$ are first- and second-order moment estimates of the gradient [49, 61]. Denoting the model weights at initialization by $\theta_{0}$, and the adaptive gradient by $\hat{g}_{t}=\hat{m}_{t}/(\sqrt{\hat{v}_{t}}+\epsilon)$, we can iterate (1) and (2) to obtain
$\theta_{T}=w_{0}^{T}\,\theta_{0}-\sum_{t=1}^{T}w_{t}^{T}\,\gamma_{t}\,\,\,\hat{g}_{t}$ (3)
where the cumulative weight decay is computed as follows
$w_{t_{1}}^{t_{2}}=\prod_{i=t_{1}+1}^{t_{2}}(1-\gamma_{i}\lambda).$ (4)
Equation (3) shows that the model weights after $t$ gradient steps are a weighted average of the initial model weights and all the adaptive gradient updates up to time step $t$. The weights $w_{t_{1}}^{t_{2}}$ determine the influence of the gradient update at time $t_{1}$ on the model weights at time $t_{2}$. With weight decay, $w_{t_{1}}^{t_{2}}$ decays to zero, as the following proposition describes.
Proposition 1.
(Cumulative Weight Decay) Fix $\epsilon>0$. Let $T=t_{2}-t_{1}$ be the number of gradient steps that have passed since step $t_{1}$. Let $\gamma_{\text{avg}}=\frac{1}{T}\sum_{t_{1}+1}^{t_{2}}\gamma_{i}$ be the average learning rate between gradient steps $t_{1}$ and $t_{2}$. If $T$ is sufficiently large, that is $T\geq\frac{\log(1/\epsilon)}{\lambda\gamma_{\text{avg}}}$, then
$w_{t_{1}}^{t_{2}}\leq\epsilon.$
We present a proof in Supplement A.5. What is the relationship between cumulative weight decay and forgetting? The adaptive gradient $\hat{g}_{t_{1}}$ contains the relevant information about the training examples seen at time $t_{1}$. If $\hat{g}_{t_{1}}$ is no longer part of the model weights because $w^{t_{2}}_{t_{1}}$ has decayed to zero, then the model should no longer depend on the examples seen at time $t_{1}$. In other words, the training examples seen at time $t_{1}$ should be forgotten. However, the data seen at time $t_{1}$ still influences $\theta_{T}$ via the trajectory of gradient descent, even if the direct contribution of $\hat{g}_{t_{1}}$ has decayed to zero. Because of this, we now investigate the relationship between forgetting and weight decay experimentally.
5.2 Empirical Forgetting Occurs Faster than the Cumulative Weight Decay
We now study the causal effect of weight decay on forgetting by intervening on the weight decay parameter, keeping everything else constant. We consider the contaminated model from Section 4.2 after two epochs and continue training with four different choices of the weight decay parameter. Figure 4 depicts the result of this experiment. Every plot in Figure 4 depicts two curves. The cumulative weight decay is depicted as a solid line. The empirically observed forgetting for 144 times repeated contamination is depicted as a dashed line (compare Figure 2(a)). From Figure 4, we see that increasing the weight decay parameter increases the empirical rate of forgetting. This is evident from the very different x-axis scales in the different plots. Moreover, the dashed line always lies below the solid line, meaning that empirical forgetting occurs faster than the cumulative weight decay. This is despite the fact that weight decay is not necessary for forgetting (see Supplement Figure 8).
@@ 5.2 Empirical Forgetting Occurs Faster than the Cumulative Weight Decay
Figure 5: Cumulative weight decay and approximate model weight composition for different large-scale training runs. Top Row: The cumulative weight decay $w_{t_{1}}^{t_{2}}$ as defined in equation (4) for different training runs. The figures depict the decay of the gradient updates for every decile of the training run, indicated in different colors. Bottom Row: The approximate composition of the final model weights in terms of the gradient updates from different deciles of the training run. The deciles are indicated in colors depicted in the legend below the plot.
5.3 Cumulative Weight Decay in Large-Scale Training
Given the discussion in Section 5.1 and the experimental results in Section 5.2, we hypothesize that example forgetting occurs at least as fast as the cumulative weight decay. Based on this hypothesis, we now gauge the degree of forgetting in large-scale training runs (via the cumulative weight decay). The cumulative weight decay only depends on the optimization hyperparameters and can be cheaply computed. [^4] Ideally, we would want experimental evidence on forgetting in large-scale training, akin to the analysis in Section 4.1. Because this exceeds our experimental budget, we rely on the assumption that forgetting in large-scale training occurs at least as fast as the cumulative weight decay, similar to what we have observed for the smaller model in Section 5.2.